If the larger of 2 nos. is divided by the smaller. The quotient and the remainder are 2 each. If 5 times the smaller no. is divided by the larger no. the quotient and the remainder are still 2 each. Find the two nos.?
Answer by edjones(8007)

Let the bigger no.=b and the smaller=s
b/s = 2 + 2/s
b=2s+2 multiply each side by s.
.
5s/b = 2 + 2/b
5s/(2s+2) = 2 + 2/(2s+2) substitute 2s+2 for b.
5s=2(2s+2)+2 multiply each side by 2s+2
5s=4s+4+2
s=6
b=14
.
Ed